OpenStudy (blackops2_143):

John’s piggy bank contained 7 dimes and 3 quarters. He pulled out 1 coin without looking. Without replacing the first coin, John then pulled out a second coin. What is the probability that both coins were dimes? Express your answer in simplest form.

Mehek (mehek14):

\(\dfrac{7}{10}*\dfrac{6}{9}=?\)

OpenStudy (blackops2_143):

56/90?

OpenStudy (blackops2_143):

or do I cross multiply it?

Mehek (mehek14):

Can it be simplified?

Mehek (mehek14):

Try dividing both by 2 56/2 90/2

OpenStudy (blackops2_143):

28 and 45

Mehek (mehek14):

so the final answer is \(\dfrac{28}{45}\)

OpenStudy (blackops2_143):

thank you!! btw I love your profile picture

Mehek (mehek14):

yw and thanks!! :)